OpenStudy (anonymous):

What is the range of the graph of y = –3(x – 2)^2 + 3? Answer A. y greater than or equal to 3 B. y greater than or equal to 2 C. y less than or equal to 3 D. y less than or equal to 2 What is the range of the graph of y = –3(x – 2)^2 + 3? Answer A. y greater than or equal to 3 B. y greater than or equal to 2 C. y less than or equal to 3 D. y less than or equal to 2 @Mathematics

OpenStudy (anonymous):

first part \[-3(x-2)^2\leq 0\] because \[(x-2)^2\geq 0\] for all x

OpenStudy (anonymous):

therefore \[-3(x-2)^2+3\leq 3\] for all x

OpenStudy (anonymous):

Thanks, Buddy :D